OpenStudy (kaylak12345):

If a cylinder has a height that is 2 times as large as its radius, and the lateral area of the cylinder is 16pi square units, then what is the length of the radius of the cylinder?

jhonyy9 (jhonyy9):

formula for lateral area of cylinder is = 2pir*h so in this case you know that the lateral area is 16pi and that h=2r , this mean that you can writing 16 = 2pir*2r 16=4pir^2 divide both sides by 4 4=pir^2 divide both sides by pi 4/pi = r^2 1,27 =r^2 r=sqrt(1,27) r= 1,12